zelma is eighteen years older than her son. she was three times as old as he was one year ago. how old are they now?
Answer by ptaylor(1894)

Let x=her son's age now
then x-1=her son's age one year ago
x+18=Zelma's age now
and x+17=Zelma's age 1 year ago
Now we are told the following:
------she was three times as old as he was (her son) one year ago-----
x+17=3(x-1) get rid of parens (distributive)
x+17=3x-3 subtract 3x and also 17 from each side
x+17-17-3x=3x-3x-3-17 collect like terms
-2x=-20 divide each side by -2
x=10-------------------------------------her son's age now
x+18=10+18=28---------------------------------Zelma's age now
CK
-----------zelma is eighteen years older than her son-----
10+18=28
28=28
------she was three times as old as he was (her son) one year ago-----
10+17=3(10-1)
27=3*9
27=27


Hope this helps---ptaylor